soybean meal is 14% protein, cornmeal is 7% protein how many pounds of each would be needed to make a 280lb mixture that is 10% protein?
Answer by checkley77(12844)

.14X+.07(280-X)=.10*280
.14X+19.6-.07X=28
.07X=28-19.6
.07X=8.4
X=8.4/.07
X=120 POUNDS OF SOYBEANS IS USED.
280-120=160 POUNDS OF CORNMEAL IS USED.
PROOF:
.14*120+.07*160=.10*280
16.8+11.2=28
28=28